### v1

# Optimal Strategy Revision in Population Games:

Abstract
This paper investigates the design of optimal strategy revision in Population Games (PG) by establishing its connection to finite-state Mean Field Games (MFG). Specifically, by linking Evolutionary Dynamics (ED) – which models agent decision-making in PG – to the MFG framework, we demonstrate that optimal strategy revision can be derived by solving the forward Fokker-Planck (FP) equation and the backward Hamilton-Jacobi (HJ) equation, both central components of the MFG framework. Furthermore, we show that the resulting optimal strategy revision satisfies two key properties: positive correlation and Nash stationarity, which are essential for ensuring convergence to the Nash equilibrium. This convergence is then rigorously analyzed and established. Additionally, we discuss how different design objectives for the optimal strategy revision can recover existing ED models previously reported in the PG literature. Numerical examples are provided to illustrate the effectiveness and improved convergence properties of the optimal strategy revision design.
Index Terms:
Population Games, Mean Field Games, Multi-Agent Strategic Interactions
I Introduction
Evolutionary Dynamics (ED) are decision-making models that guide a large population of agents to learn strategies that form a Nash equilibrium in Population Games (PG). These models are characterized by strategy revision protocols, which define how agents revise between strategies based on the payoffs received from their current strategy choices. While various revision protocols have been studied, a key question remains unanswered: which protocol is the most effective for achieving the equilibrium, given a specified objective?
Recent advances in the PG literature have introduced higher-order dynamics models for refining payoff determinations [1, 2]. These models, for example, can be designed to filter out noise in perceived payoffs [3] or to mitigate the adverse effects of time delays on the convergence to the Nash equilibrium [4]. In this work, leveraging the Mean Field Games (MFG) framework, we propose a novel approach to designing a payoff dynamics model that optimally balances rapid convergence to the Nash equilibrium with the penalties incurred during strategy revision.
Solutions for standard MFG problems are characterized by two equations: a backward Hamilton-Jacobi equation, which describes the evolution of the optimal cost-to-go, and a forward Fokker-Planck-Kolmogorov equation, which captures the distribution of agents’ states. Although both MFG and PG focus on decision-making in large populations, their primary difference lies in the structure of their strategy sets: MFG typically involves a continuous strategy set, as seen in differential games of the mean-field type [5], whereas PG is defined over a finite number of strategies [6].
The study by [7] introduces a finite-state MFG framework, laying the groundwork for establishing a connection between MFG and PG. Building on this foundation, the primary contribution of our work is to formally establish the relationship between MFG and PG frameworks. Specifically, we demonstrate the connection between the Fokker-Planck (FP) equation in finite-state MFG and the ED model in PG. Leveraging this connection, we show that the optimal solution in the finite-state MFG framework corresponds to an optimal strategy revision protocol in PG, with payoff dynamics derived from the backward Hamilton-Jacobi (HJ) equation. To the best of our knowledge, this is the first work to formally establish a direct connection between MFG and ED models in PG.
The contributions of this paper can be summarized as follows: $i)$ We establish a connection between the finite-state MFG introduced in [7] and the ED models presented in [6, 8]. This is achieved by introducing a dynamic model for payoff vectors analogous to the backward HJ equation in MFG. $ii)$ We demonstrate that, under the proposed payoff dynamics model, changes in the distribution of the agents’ strategy selections exhibit positive correlation with the payoff vector. This behavior aligns with the positive correlation property observed in conventional ED models and is critical for ensuring convergence to the Nash equilibrium. Additionally, we establish the link between the steady-state solution of the MFG model and the Nash equilibrium in PG. The convergence of the population state to the Nash equilibrium is rigorously discussed. $iii)$ We show that existing ED models, including Smith, replicator, and projection dynamics, can be derived as special cases of our approach.
This paper is organized as follows: §$\S$§II provides the problem statement and a brief overview of preliminaries on PG and finite-state MFG. §$\S$§III presents the main results of this paper. Illustrative examples are discussed in §$\S$§IV. Finally, conclusions are drawn in §$\S$§V.
II Preliminaries and Problem Statement
II-A Population Games and Problem Formulation
Consider a large population of decision-making agents, each selecting a strategy from the set $\mathbb{S}=\{1,\cdots,n\}$. At any time $t\in[t_{0},T]$ with $t_{0}<T$, agents may revise their strategy choices. The vector-valued function $x:[t_{0},T]\to\Delta$ represents the state of the population, describing the distribution of agents’ strategy selections at a given time $t$. The set $\Delta$, known as the probability simplex, is defined as $\Delta=\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{i\in\mathbb{S}}x_{i}=1\}$, representing the set of all feasible population states. The $i$-th entry $x_{i}(t)$ of $x(t)$ specifies the proportion of agents selecting strategy $i\in\mathbb{S}$ at time $t$.
In standard population games, the payoff function $F:\Delta\to\mathbb{R}^{n}$ specifies the payoff vector $F(x(t))$ associated with the population state $x(t)$. For a given $F$, the Nash equilibrium is defined as follows.
Definition 1
A population state $x^{*}\in\Delta$ is called a Nash equilibrium of $F$ if the following holds for all $i$ in $\mathbb{S}$:
$x_{i}^{*}>0\implies F_{i}(x^{*})=\textstyle\max_{j\in\mathbb{S}}F_{j}(x^{*}).$
The set of all Nash equilibria is denoted by $\mathbb{NE}(F)$.
The definition of contractive population games is given as follows.
Definition 2
A population game $F$ is called contractive if the following inequality holds for all $x,y$ in $\Delta$:
$\left(F(x)-F(y)\right)^{\top}\left(x-y\right)\leq 0.$ (1)
Agents are allowed to revise their strategy selections at the arrival times of i.i.d. Poisson processes [6]. Let $\rho_{ji}(p,x)$ denote the strategy revision protocol, which specifies the likelihood that a strategy-revising agent switches from strategy $j$ to strategy $i$. For instance, the Smith protocol is defined as $\rho_{ji}(p,x)=[p_{i}-p_{j}]_{+}$, where $[\cdot]_{+}=\max(0,\cdot)$. As shown in [6], under the assumptions of a homogeneous population – where all agents adopt the same revision protocol – and an infinitely large population, the population state $x(t)$ evolves according to the following ordinary differential equation, commonly referred to as the Evolutionary Dynamics (ED):
$\dot{x}_{i}(t)=\textstyle\sum_{j\in\mathbb{S}}x_{j}(t)\rho_{ji}(p(t),x(t))\\ -x_{i}(t)\textstyle\sum_{j\in\mathbb{S}}\rho_{ij}(p(t),x(t)),~{}i\in\mathbb{S},$ (2)
where $p(t)=F(x(t))$. Let $V$ be the vector field defining $\dot{x}(t)=V(p(t),x(t))$, where $x(t)=(x_{1}(t),\cdots,x_{n}(t))$. As discussed in [9], the following key properties of ED are crucial in establishing convergence to the Nash equilibrium.
Definition 3
ED satisfies $i)$ the Nash stationarity if $V(F(x),x)=0$ holds if and only if $x\in\mathbb{NE}(F)$, and $ii)$ the positive correlation if $V(p,x)\neq\mathbf{0}$ implies $p^{\top}V(p,x)>0$.
The primary objective of this work is to design an optimal strategy revision protocol $\rho=(\rho_{ij})_{i,j\in\mathbb{S}}$ and a payoff dynamics model
$\dot{p}_{i}(t)=G_{i}(t,p(t),x(t)),~{}i\in\mathbb{S},$ (3)
that maximize the agents’ long-term payoffs. The problem is formulated as an optimal control problem for a reference agent starting with each strategy $i$. The reference agent is defined as a randomly selected agent whose strategy selection at time $t_{0}$ is $i$ [^1] In population games, agents are non-atomic, meaning their individual identities are negligible; only their strategy choices matter. Consequently, the finite time-horizon payoff function (4) describes the long-term payoffs received by agents who begin with strategy $i$ under the revision protocol $\rho$.
Main Problem: Consider an optimal control problem where a reference agent initially selects strategy $i$ at time $t_{0}$ and subsequently revises its strategy according to the strategy revision protocol. The objective is to design $\rho=(\rho_{ij})_{i,j\in\mathbb{S}}$ and $G=(G_{i})_{i\in\mathbb{S}}$ to maximize a finite time-horizon payoff function defined as: [^2] The expectation in (4) is taken with respect to the strategy selection $s(t)$, a random variable over $\mathbb{S}$ with the transition probability satisfying [10, §4]: $\lim_{\epsilon\to 0}\frac{1}{\epsilon}\mathbb{P}\left[s(t+\varepsilon)=j\,\big{|}\,s(t)=i\right]=\rho_{ij}(p(t),x(t))$.
$P_{i}(t_{0},\rho)=\mathbb{E}\Big{[}\textstyle\int_{t_{0}}^{T}\big{(}-\sum_{j\in\mathbb{S}}\frac{q_{s(t),j}(t)}{2}\rho_{s(t),j}^{2}(p(t),x(t))\\ +F_{s(t)}(x(t))\big{)}\mathrm{d}t+F_{s(T)}(x(T))\,\Big{|}\,s(t_{0})=i\Big{]},$ (4)
where $F=(F_{1},\cdots,F_{n})$ is the instantaneous payoff function of the underlying population game, $s(t)\in\mathbb{S}$ denotes the agent’s strategy at time $t$, and $x(t)$ and $p(t)$ are determined by (2) and (3), respectively.
The finite time-horizon payoff function $P_{i}(t_{0},\rho)$ accumulates payoffs over $[t_{0},T]$, including a penalty term proportional to the square of the protocol $\rho_{ji}$. The weight $q_{ij}(t),~{}i,j\in\mathbb{S}$ plays a crucial role in the optimal control problem. Reducing $q_{ij}(t)$ decreases the penalization for revising strategies from $i$ to $j$, thereby accelerating the dynamics in (2). The weight $q_{ij}(t)$ may depend on the population states, as discussed in §$\S$§III-A.
II-B Finite-State Mean-Field Games
Consider a large population of homogeneous agents interacting within a finite-state MFG framework over a time horizon $[t_{0},T]$ with $t_{0}<T$, as discussed in [7]. In this framework, the state of a reference agent is modeled by a continuous-time Markov chain defined over a finite state space $\mathbb{S}=\{1,\cdots,n\}$. State transitions are governed by an $n\times n$ transition matrix $\alpha(t)=(\alpha_{ij}(t))_{i,j\in\mathbb{S}}\in\mathbb{R}^{n\times n}$, satisfying the following conditions at each time $t$: $\alpha_{ij}(t)\geq 0,~{}\forall i,j\in\mathbb{S}$, $i\neq j$, and $\sum_{j\in\mathbb{S}}\alpha_{ij}(t)=0,~{}\forall i\in\mathbb{S}$. More formally, for $\varepsilon>0$, the transition probability is defined as
$\mathbb{P}\left[s(t+\varepsilon)=j\,\big{|}\,s(t)=i\right]=\alpha_{ij}(t)\varepsilon+o(\varepsilon),$ (5)
where $s(t)\in\mathbb{S}$ represents the agent’s state at time $t$, and $\lim_{\varepsilon\to 0}o(\varepsilon)/\varepsilon=0$.
The state distribution across $\mathbb{S}$ is represented by a function $x:[t_{0},T]\to\Delta$, referred to as the mean-field term. The main objective in MFG problems is to compute the transition matrix $\alpha(t),\,t\in[t_{0},T]$ that minimizes a finite time-horizon cost function $J_{i}$ for a reference agent starting in state $i$ at time $t_{0}$, defined as:
$J_{i}(t_{0},\alpha)=\mathbb{E}\Big{[}\textstyle\int_{t_{0}}^{T}c(s(t),x(t),\alpha_{s(t)}(t))\,\mathrm{d}t\\ +\psi(s(T),x(T))\,\Big{|}\,s(t_{0})=i\Big{]},$ (6)
where $\alpha_{i}(t)=(\alpha_{i1}(t),\cdots,\alpha_{in}(t))$ is the $i$-th row of the transition matrix, $c:\mathbb{S}\times\Delta\times\mathbb{R}^{n}\to\mathbb{R}$ is the running cost, and $\psi:\mathbb{S}\times\Delta\to\mathbb{R}$ is the terminal cost.
For notational convenience, given $i$ in $\mathbb{S}$, let $\mathbb{R}_{i,0}^{n}$ denote the set of vectors $\mu=(\mu_{1},\cdots,\mu_{n})\in\mathbb{R}_{i,0}^{n}$ satisfying: $i)$ $\mu_{j}\geq 0$ if $j\neq i$, and $ii)$ $\sum_{j\in\mathbb{S}}\mu_{j}=0$. The optimal cost-to-go from time $t$ is defined as $w_{i}(t,x)=\min_{\alpha}J_{i}(t,\alpha)$.
The solution to the MFG problem is given by two coupled equations, as outlined in [7]. The first is the backward evolution of $w$, known as the Hamilton-Jacobi (HJ) equation:
$\dot{w}_{i}(t,x(t))=-H_{i}(w(t,x(t)),x(t))$ (7a)
$w_{i}(T,x(T))=\psi(i,x(T)),$ (7b)
where the Hamiltonian $H_{i}$ is defined as
$H_{i}(z,x)=\textstyle\min_{\mu\in\mathbb{R}_{i,0}^{n}}\left(c(i,x,\mu)+\sum_{j\in\mathbb{S}}\mu_{j}(z_{j}-z_{i})\right).$
The second equation describes the forward evolution of $x(t)$, commonly referred to as the Fokker-Plank (FP) equation:
$\begin{array}[]{l}\dot{x}_{i}(t)=\textstyle\sum_{j\in\mathbb{S}}\alpha_{ji}(t)x_{j}(t),~{}x_{i}(t_{0})=x_{0,i},\end{array}$ (9)
where $(x_{0,1},\cdots,x_{0,n})\in\Delta$, and
$\alpha_{i}(t)=\textstyle\operatorname*{arg\,min}_{\mu\in\mathbb{R}_{i,0}^{n}}\Big{(}c(i,x(t),\mu)\\ +\sum_{j\in\mathbb{S}}\mu_{j}(w_{j}(t,x(t))-w_{i}(t,x(t)))\Big{)},$
representing the optimal state transition. For the existence of solutions to (7) and (9), see [7, Proposition 4].
The MFG system (7) and (9) may not always admit a stationary solution, defined as $\dot{w}(t,x(t))=\dot{x}(t)=0$, for instance, when $H_{i}>0$. In such cases, we define a stationary solution $(\bar{w},\bar{x})\in\mathbb{R}^{n}\times\Delta$ (modulo addition of a constant $\kappa$) as follows: For all $i$ in $\mathbb{S}$,
$H_{i}(\bar{w},\bar{x})$ $=\kappa$ (10a)
$\textstyle\sum_{j\in\mathbb{S}}\bar{\alpha}_{ji}\bar{x}_{j}$ $=0,$ (10b)
where $\bar{\alpha}_{i}=\operatorname*{arg\,min}_{\mu\in\mathbb{R}_{i,0}^{n}}(c(i,\bar{x},\mu)+\sum_{j\in\mathbb{S}}\mu_{j}(\bar{w}_{j}-\bar{w}_{i}))$.
III Optimal Strategy Revision Protocol Design
Figure 1: Closed-loop model diagrams for (a) MFG and (b) the payoff dynamics for the optimal strategy revision in PG. To highlight the relationship, recall that $v_{i}(t,x(t))=-w_{i}(t,x(t))$.
We begin by establishing the relationship between the MFG system and the optimal strategy revision process in PG, as illustrated in Fig. 1. Lemma 1 shows the connection between the state distribution evolution in finite-state MFG and the ED model, followed by Theorem 1, which provides a solution to our main problem.
Lemma 1
The ED (2) is equivalent to the FK equation (9) if the strategy revision protocol satisfies
$\rho_{ij}(p(t),x(t))=\begin{cases}\alpha_{ij}(t)&\text{if}~{}i\neq j\\ 0&\text{otherwise}\end{cases},~{}i,j\in\mathbb{S}.$ (11)
Proof:
By definition, we have $\sum_{j\in\mathbb{S}\setminus\{i\}}\alpha_{ij}(t)=-\alpha_{ii}(t)$, which allows us to rewrite (9) as:
$\dot{x}_{i}(t)$ $=\textstyle\sum_{j\in\mathbb{S}\setminus\{i\}}\alpha_{ji}(t)x_{j}(t)+\alpha_{ii}(t)x_{i}(t).$
This simplifies further to:
$\dot{x}_{i}(t)\!=\!\textstyle\sum_{j\in\mathbb{S}\setminus\{i\}}x_{j}(t)\alpha_{ji}(t)\!-\!x_{i}(t)\sum_{j\in\mathbb{S}\setminus\{i\}}\alpha_{ij}(t).$ (12)
Notably, the value of $\alpha_{ii}(t)$ does not influence (12) because the sums can be extended to include all $j\in\mathbb{S}$ without altering the results, as the contributions for $j=i$ are null. With $\rho_{ij}(p(t),x(t))=\alpha_{ij}(t)$ for all $i\neq j$, we complete the proof. [^3] Also, in the context of ED, when considering pairwise comparison revision protocols, it holds that $\rho_{ii}(p(t),x(t))=0$, for all $i\in\mathbb{S}$, showing that (2) can be written using the same sum indexing as in (12), i.e., over $j\in\mathbb{S}\setminus\{i\}$.   $\blacksquare$
Theorem 1
The optimal strategy revision protocol for (4) is given by:
$\rho_{ji}(p(t),x(t))=\frac{[p_{i}(t)-p_{j}(t)]_{+}}{q_{ji}(t)},$
where $p_{i}(t)=v_{i}(t,x(t))$, and $v_{i}(t,x(t))$ satisfies the following backward differential equation:
$\dot{v}_{i}(t,x(t))$ $=-\frac{1}{2}\textstyle\sum_{j\in\mathbb{S}}\dfrac{[v_{j}(t,x(t))\!-\!v_{i}(t,x(t))]_{+}^{2}}{q_{ij}(t)}$
$\qquad-F_{i}(x(t))$ (13a)
$v_{i}(T,x(T))$ $=F_{i}(x(T)).$ (13b)
The corresponding evolution of $x_{i}(t)$ is described by:
$\dot{x}_{i}(t)$ $=\textstyle\sum_{j\in\mathbb{S}}x_{j}(t)\dfrac{[v_{i}(t,x(t))-v_{j}(t,x(t))]_{+}}{q_{ji}(t)}$
$\quad-x_{i}(t)\textstyle\sum_{j\in\mathbb{S}}\dfrac{[v_{j}(t,x(t))-v_{i}(t,x(t))]_{+}}{q_{ij}(t)},$ (14a)
$x_{i}(t_{0})$ $=x_{0,i},$ (14b)
where $x({t_{0}})\in\Delta$.
Proof:
We begin by determining the optimal state transition matrix $\alpha(t)$ within the MFG framework, where the finite time-horizon cost function in (6) is defined with $c(s(t),x(t),\alpha_{s(t)}(t))=\sum_{j\in\mathbb{S}}\frac{q_{s(t),j}(t)}{2}\alpha_{s(t),j}^{2}(t)-F_{s(t)}(x(t))$ and $\psi(s(T),x(T))=-F_{s(T)}(x(T))$. The Hamiltonian is given by:
$H_{i}(w(t,x(t)),x(t))=\textstyle\min_{\alpha_{i}\in\mathbb{R}_{i,0}^{n}}\bigg{(}\sum_{j\in\mathbb{S}}\frac{q_{ij}(t)}{2}\alpha_{ij}^{2}\\ -F_{i}(x(t))+\textstyle\sum_{j\in\mathbb{S}}\alpha_{ij}(w_{j}(t,x(t))-w_{i}(t,x(t)))\bigg{)}.$
Under the constraint that $\alpha_{i}$ must lie in $\mathbb{R}_{i,0}^{n}$, the optimal state transition $\alpha_{ij}(t)$ is then given by:
$\alpha_{ij}(t)=\frac{[w_{i}(t,x(t))-w_{j}(t,x(t))]_{+}}{q_{ij}(t)},~{}\forall i\neq j.$ (15)
Consequently, we derive the Hamiltonian function as follows:
$H_{i}(w(t,x(t)),x(t))\\ =-\frac{1}{2}\textstyle\sum_{j\in\mathbb{S}}\dfrac{[w_{i}(t,x(t))-w_{j}(t,x(t))]_{+}^{2}}{q_{ij}(t)}-F_{i}(x(t)),$
for all $i$ in $\mathbb{S}$. Since $w_{i}(t,x)$ is the value function of (6), we define $v_{i}(t,x)=-w_{i}(t,x)$, making $v_{i}(t,x)$ the optimal cost-to-go for (4). By combining (7) with Lemma 1, we use $\alpha_{ij}(t)$ and $v_{i}(t,x(t))$ to define $\rho_{ij}(p(t),x(t))$ and $p_{i}(t)$, respectively. This concludes the proof. $\blacksquare$
As stated in Theorem 1, the payoff dynamics model (3) is given by:
$G_{i}(t,p(t),x(t))=-\frac{1}{2}\textstyle\sum_{j\in\mathbb{S}}\dfrac{[p_{j}(t)-p_{i}(t)]_{+}^{2}}{q_{ij}(t)}-F_{i}(x(t))$
with the terminal condition $p_{i}(T)=F_{i}(x(T))$ for all $i$ in $\mathbb{S}$.
III-A Direct Connection with Conventional ED Models
The following corollary demonstrates how specific classes of existing ED models can be derived from Theorem 1 by identifying a corresponding weight $q_{ij}(t)$ in (4) for each class. Each resulting ED model is named after its counterpart in the literature, as detailed in Remark 1, where we provide an interpretation of the associated revision protocol. This interpretation offers valuable insights into the behaviors observed in PG.
Corollary 1
Consider the following choices for the weight $q_{ij}(t)$ in (4): $i)$ $q_{ij}(t)=1$, $ii)$ $q_{ij}(t)=1/x_{j}(t)$, and $iii)$ $q_{ij}(t)=x_{i}(t)$. For each case, the corresponding ED models are as follows:
$i)$ Smith dynamics and payoff dynamics:
$\dot{x}_{i}(t)$ $=\textstyle\sum_{j\in\mathbb{S}}x_{j}(t)[v_{i}(t,x(t))-v_{j}(t,x(t))]_{+}$
$\qquad-x_{i}(t)\textstyle\sum_{j\in\mathbb{S}}[v_{j}(t,x(t))-v_{i}(t,x(t))]_{+}$
$\dot{v}_{i}(t,\!x(t))$ $\!=\!-\frac{1}{2}\textstyle\sum_{j\in\mathbb{S}}[v_{j}(t,\!x(t))\!-\!v_{i}(t,\!x(t))]_{+}^{2}\!-\!F_{i}(x(t)).$
$ii)$ Replicator dynamics and payoff dynamics:
$\dot{x}_{i}(t)$ $=x_{i}(t)\big{(}v_{i}(t,x(t))-\textstyle\sum_{j\in\mathbb{S}}x_{j}(t)v_{j}(t,x(t))\big{)}$
$\dot{v}_{i}(t,x(t))$ $=-\frac{1}{2}\textstyle\sum_{j\in\mathbb{S}}x_{j}(t)[v_{j}(t,x(t))-v_{i}(t,x(t))]_{+}^{2}$
$\qquad-F_{i}(x(t)).$
$iii)$ Projection dynamics and payoff dynamics:
$\dot{x}_{i}(t)$ $=\textstyle\sum_{j\in\mathbb{S}}\big{(}v_{i}(t,x(t))-v_{j}(t,x(t))\big{)}$
$\dot{v}_{i}(t,\!x(t))$ $\!=\!-\frac{1}{2}\textstyle\sum_{j\in\mathbb{S}}\dfrac{[v_{j}(t,\!x(t))\!-\!v_{i}(t,x(t))]_{+}^{2}}{x_{i}(t)}-F_{i}(x(t)).$
All models satisfy the terminal condition $v_{i}(T,x(T))=F_{i}(x(T))$ and the initial condition $x(t_{0})\in\Delta$.
Corollary 1 highlights interesting connections between the choice of $q_{ij}(t)$ and several well-known properties in ED, as elaborated in the following remarks.
Remark 1
Note the following two facts:
$i)$ For the replicator dynamics, (4) diverges as the proportion of decision-makers approaches zero, i.e., $q_{ij}(t)\to\infty$ when $x_{j}(t)\to 0$. Consequently, choosing $q_{ij}(t)={1}/{x_{j}(t)}$ acts as a penalty function, ensuring that the boundary $\partial\Delta$ of the simplex is never reached. This aligns with the observation that under the replicator dynamics, the boundary $\partial\Delta$ acts as a repeller. This interpretation is consistent with the permanence theorem discussed in [11].
$ii)$ As $T\to t_{0}$ in (13), we recover $v_{i}(t_{0},x(t_{0}))=F_{i}(x(t_{0}))$. Substituting this result into (14) reproduces the standard ED models, as in (2), widely studied in the literature [6, 12, 13].
In the following corollary, we consider migration constraints that restrict agents from switching between specific pairs of strategies. These permissible strategy switches are represented by an undirected connected graph $(\mathbb{S},\mathbb{L},A)$, where $\mathbb{L}\subseteq\{(i,j)\,|\,i,j\in\mathbb{S}\}$ denotes the set of allowed migrations (graph links), and a symmetric matrix $A\in\{0,1\}^{n\times n}$ is the adjacency matrix. The entries $a_{ij}$ of $A$ capture the migration constraints, such that $(i,j)\in\mathbb{L}$ if and only if $a_{ij}=1$. The set of neighbors of strategy $i$ is defined as $\mathbb{S}_{i}=\{j\in\mathbb{S}\,|\,(j,i)\in\mathbb{L}\}$. Prior works [12, 13, 14] have demonstrated that migration constraints lead to distributed population dynamics by incorporating the adjacency matrix $A$ – or equivalently, the migration constraints – into the revision protocol.
Corollary 2
Consider the following choices for the weights $q_{ij}(t)$ in (4) using $A=(a_{ij})_{i,j\in\mathbb{S}}$: $i)$ $q_{ij}(t)={1}/{a_{ij}}$, $ii)$ $q_{ij}(t)={1}/{(a_{ij}x_{j}(t))}$, and $iii)$ $q_{ij}(t)={x_{i}(t)}/{a_{ij}}$. For each case, the corresponding ED models are as follows:
$i)$ Distributed Smith dynamics and payoff dynamics:
$\dot{x}_{i}(t)$ $=\textstyle\sum_{j\in\mathbb{S}_{i}}x_{j}(t)[v_{i}(t,x(t))-v_{j}(t,x(t))]_{+}$
$\qquad-x_{i}(t)\textstyle\sum_{j\in\mathbb{S}_{i}}[v_{j}(t,x(t))\!-\!v_{i}(t,x(t))]_{+}$
$\dot{v}_{i}(t,x(t))$ $=-\frac{1}{2}\textstyle\sum_{j\in\mathbb{S}_{i}}[v_{j}(t,x(t))-v_{i}(t,x(t))]_{+}^{2}\!-\!F_{i}(x(t)).$
$ii)$ Distributed replicator dynamics and payoff dynamics:
$\dot{x}_{i}(t)$ $=x_{i}(t)\big{(}v_{i}(t,x(t))\textstyle\sum_{j\in\mathbb{S}_{i}}x_{j}(t)$
$\qquad\qquad-\textstyle\sum_{j\in\mathbb{S}_{i}}x_{j}(t)v_{j}(t,x(t))\big{)}$
$\dot{v}_{i}(t,x(t))$ $=-\frac{1}{2}\textstyle\sum_{j\in\mathbb{S}_{i}}x_{j}(t)[v_{j}(t,x(t))-v_{i}(t,x(t))]_{+}^{2}$
$\qquad-F_{i}(x(t)).$
$iii)$ Distributed projection dynamics and payoff dynamics:
$\dot{x}_{i}(t)$ $=\textstyle\sum_{j\in\mathbb{S}_{i}}(v_{i}(t,x(t))-v_{j}(t,x(t)))$
$\dot{v}_{i}(t,x(t))$ $\!=\!-\frac{1}{2}\textstyle\sum_{j\in\mathbb{S}_{i}}\dfrac{[v_{j}(t,x(t))\!-\!v_{i}(t,x(t))]_{+}^{2}}{x_{i}(t)}\!-\!F_{i}(x(t)).$
All models satisfy the terminal condition $v_{i}(T,x(T))=F_{i}(x(T))$ and the initial condition $x(t_{0})\in\Delta$.
III-B Positive Correlation, Nash Stationarity, and Convergence
In the next two propositions, we show that the optimal strategy revision protocol satisfies positive correlation and Nash stationarity, similar to standard ED models (see Definition 3)
Proposition 1
The payoff dynamics model (13) and the ED model (14) satisfy the following positive correlation: $V(p(t),x(t))\neq\mathbf{0}$ implies $p^{\top}(t)V(p(t),x(t))>0$, where $p(t)=v(t,x(t))$, for all $t$ in $[t_{0},T]$.
Proof:
Using (14), we can evaluate $p^{\top}(t)V(p(t),x(t))$ as follows:
$p^{\top}(t)V(p(t),x(t))$
$=\textstyle\sum_{i\in\mathbb{S}}v_{i}(t,x(t))\dot{x}_{i}(t)$
$=\textstyle\sum_{i\in\mathbb{S}}\sum_{j\in\mathbb{S}}x_{j}(t)\dfrac{[v_{i}(t,x(t))-v_{j}(t,x(t))]_{+}^{2}}{q_{ji}(t)}.$
Therefore, we conclude that $p^{\top}(t)V(p(t),x(t))\geq 0$, with equality holds only if $V(p(t),x(t))=0$. This concludes the proof. $\blacksquare$
Proposition 1 implies that, under the optimal strategy revision protocol $\rho$, the population state evolves to maximize the finite time-horizon payoff function (4).
Proposition 2
Consider the modulo-addition stationary solution of the associated MFG system, as explained in (10). The stationary solution for (13) and (14) is given by: $v(t,\bar{x})=\kappa(t-t_{0})\mathbf{1}_{n}+v(t_{0},\bar{x}),~{}\forall t\in[t_{0},T]$, where $\bar{x}$ is a Nash equilibrium of $F$, and $v(T,\bar{x})$ is its associated payoff vector given by $v(T,\bar{x})=F(\bar{x})$.
Proof:
From (10a), we have:
$\dot{v}(t,\bar{x})=\kappa\mathbf{1}_{n}\implies v(t,\bar{x})=\kappa(t-t_{0})\mathbf{1}_{n}+v(t_{0},\bar{x}),$
with $v(T,\bar{x})=F(\bar{x})$. From (10b), it follows that $V(p(t),\bar{x})=\mathbf{0}$. Following the same reasoning as in the proof of Proposition 1, this condition implies $\bar{x}_{i}[F_{j}(\bar{x})-F_{i}(\bar{x})]_{+}=0,~{}\forall i,j\in\mathbb{S}$, thereby establishing that $\bar{x}$ is a Nash equilibrium.  $\blacksquare$
In the following corollary, we provide a convergence result for a certain class of contractive population games.
Corollary 3
Suppose a population game $F$ is contractive and satisfies
$\left(F(x)-F(y)\right)^{\top}(x-y)\leq-\epsilon\|x-y\|_{2}^{2},$ (19)
where $\epsilon$ is a positive constant. Then, $x(t)$ converges to a Nash equilibrium of $F$.
Proof:
By Proposition 2, the modulo-addition stationary solution corresponds to a Nash equilibrium $x^{\ast}$. Invoking [7, Theorem 3], which states that $x(t)$ converges to the modulo-addition stationary solution when $F$ satisfies (19), we conclude that $x(t)$ converges to $x^{\ast}$. $\blacksquare$
Remark 2
For any contractive population game $F$, modifying $F$ by defining $F_{\epsilon}(x)=F(x)-\epsilon\,x$ with any positive $\epsilon$ ensures that $F_{\epsilon}$ satisfies (19). Let $x^{\ast}(\epsilon)$ denote the Nash equilibrium of $F_{\epsilon}$, which also serves as the $\epsilon$-Nash equilibrium of $F$. If $F$ is a continuous function of $x$, it follows that:
$\lim_{\epsilon\to 0}x^{\ast}(\epsilon)=x^{\ast},$
where $x^{\ast}$ is the Nash equilibrium of $F$.
IV Illustrative Examples
Data: $F$, $x_{0})\in\Delta$, $t_{0},T\in\mathbb{R}_{+}$ with $T>t_{0}$, $N\in\mathbb{Z}_{+}$, $\varepsilon_{f}\in\mathbb{R}_{+}$
Result: $x^{*}_{[t_{0},T]},v^{*}_{[t_{0},T]}$
1
[…]
3
[…]
5
6while $(k<N)$ and $(e>\varepsilon_{f})$ do
7
8      $v^{(k+1)}_{[t_{0},T]}\leftarrow$ a solution of the HJ equation (13) with a fixed population state trajectory $x_{[t_{0},T]}$ as $x_{[t_{0},T]}=x^{(k)}_{[t_{0},T]}$
9       $x^{\prime}_{[t_{0},T]}\leftarrow$ a solution of the FK equation (14) with a fixed payoff vector trajectory $p_{[t_{0},T]}$ as $p_{[t_{0},T]}=v^{(k+1)}_{[t_{0},T]}$
10       $x^{(k+1)}_{[t_{0},T]}\leftarrow(1-\alpha)x^{(k)}_{[t_{0},T]}+\alpha x^{\prime}_{[t_{0},T]}$
11       $e\leftarrow\int_{t_{0}}^{T}\|x^{(k)}(t^{\prime})-x^{(k+1)}(t^{\prime})\|_{2}^{2}\,\mathrm{d}t^{\prime}$triangleright To check if the fixed-point is reached
12       $k\leftarrow k+1$
13
$(x^{*}_{[t_{0},T]},v^{*}_{[t_{0},T]})\leftarrow(x^{(k+1)}_{[t_{0},T]},v^{(k+1)}_{[t_{0},T]})$
Algorithm 1 Compute solution trajectories $(x^{*}_{[t_{0},T]},v^{*}_{[t_{0},T]})$ of (13) and (14).
We provide numerical examples to illustrate the main results of this paper. To compute the solutions for (13) and (14), we use Algorithm 1, adapted from the literature [15]. The algorithm is specifically designed to find the fixed point of (13) and (14).
The computation begins by initializing the population state trajectory using a conventional game model, $p=F(x)$. The algorithm then alternates between updating the payoff vector trajectory based on the current population state trajectory and vice versa. To stabilize this alternating process, we employ an exponential moving average strategy (Line 6) when updating the population state trajectory. This approach helps mitigate rapid changes and ensures smoother convergence. Each iteration in the algorithm is denoted by the superscript $(k)$.
We consider the following two population games.
Congestion game
The payoff function $F$ is defined as:
$F(x)=\begin{pmatrix}F_{1}(x)\\ F_{2}(x)\\ F_{3}(x)\end{pmatrix}=-\begin{pmatrix}3x_{1}+x_{3}\\ 2x_{2}+x_{3}\\ x_{1}+x_{2}+3x_{3}\end{pmatrix}.$ (20)
For (4), we set $q_{ij}=1,\forall i,j\in\mathbb{S}$, $t_{0}=0$, and $T=6$. We apply Algorithm 1 with $\alpha=0.01$, $N=100$, and $\varepsilon_{f}=0$.
Rock-Paper-Scissors (RPS) game
The payoff function $F$ is defined as:
$F(x)=\begin{pmatrix}F_{1}(x)\\ F_{2}(x)\\ F_{3}(x)\end{pmatrix}=\begin{pmatrix}-x_{2}+x_{3}\\ x_{1}-x_{3}\\ -x_{1}+x_{2}\end{pmatrix}.$ (21)
For (4), we set $q_{ij}=1,\forall i,j\in\mathbb{S}$, $t_{0}=0$, and $T=6$. We apply Algorithm 1 with $\alpha=0.001$, $N=6000$, and $\varepsilon_{f}=0$.
Figure 2: (a) The progression of the error term in Algorithm 1, and (b) the resulting trajectory of the population state $x_{2}^{(k)}(t)$ over various iterations in the congestion game (20).
Figure 3: A comparison between population state trajectories determined by the optimal strategy revision protocol and the Smith protocol.
Fig. 2 illustrates the progression of the error term $\int_{t_{0}}^{T}\|x^{(k)}(t^{\prime})-x^{(k+1)}(t^{\prime})\|_{2}^{2}\,\mathrm{d}t^{\prime}$ (Line 7) and the iterative improvements of $x_{2}^{(k)}(t)$ during the execution of Algorithm 1 for the congestion game (20). As shown in Fig. 2, the overshoot in $x_{2}^{(k)}(t)$ decreases with each iteration, leading to a reduced cost of strategy revision and an increase in the long-term payoff (4).
Fig. 3 compares the trajectories generated by the optimal strategy revision protocol with those obtained using the Smith ED model in the RPS game (21). As shown, the population state trajectory determined by the optimal protocol exhibits reduced oscillations and converges more quickly to the Nash equilibrium of (21).
V Conclusions
This paper examined the relationship between the MFG framework and the optimal strategy revision process in PG. Specifically, we established a connection between the evolution of agent state distributions in MFG, governed by the forward FK equation, and the ED model in PG. Furthermore, we demonstrated that the backward HJ equation in MFG can be utilized to define the payoff dynamics model, enabling the determination of the payoff vector for the optimal strategy revision in PG. Building on these insights, we showed how existing ED models – including Smith, replicator, and projection dynamics, along with their distributed variants – can be derived by appropriately selecting the finite time-horizon payoff function.
One limitation of the proposed approach is that agents require full knowledge of the payoff function $F$ of the underlying population game. As a direction for future research, we plan to investigate learning-based approaches where agents, through repeated interactions within the same game, can learn the optimal strategy revision protocol without requiring this assumption.